# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 1.1. Introduction
$\frac{P_{1}}{\sqrt{T_{1}}}=\frac{P_{2}}{\sqrt{T_{2}}}$ (1.1)
in equilibrium. This phenomenon is called the thermomolecular pressure difference. This relation is discussed in [25], [26]. In an experiment, S. Chu Liang provided a more accurate expression of (1.1) in [27], with further applications given in [2], [40].
@@
$v\cdot\nabla_{x}F(x,v)=\frac{1}{\kappa}Q(F,F)(x,v),\quad\text{in}\quad\Omega\times\mathbb{R}^{3}.$ (1.2)
@@
The collision operator $Q$ that appears in (1.2) is defined as follows.
@@
$v^{\prime}=v+((u-v)\cdot\omega)\omega\quad\text{and}\quad u^{\prime}=u-((u-v)\cdot\omega)\omega.$ (1.3)
@@
$v^{\prime}=\frac{v+u}{2}+\frac{|v-u|}{2}\sigma\quad\text{and}\quad u^{\prime}=\frac{v+u}{2}-\frac{|v-u|}{2}\sigma,$ (1.4)
@@
Consequently, stationary Boltzmann equation has been extensively applied to describe the gas distribution in the domains involving thermal transpiration. However, in engineering applications, obtaining numerical solutions requires substantial computational efforts, either deterministically ([38], [42], [45]) or stochastically ([3], [33]). Some numerical analysis of thermal transpiration employ the lattice Boltzmann model [39], the BGK model ([20], [28], [30]) or the linearized Boltzmann Equation model [41]. In particular, for the linearized Boltzmann equation model, Sone, Y. and Aoki, K. have made many significant contributions ([KazuoAokiYoshioSoneandYorifumiWaniguchi], [31]). Moreover, Sone provided an explicit form of the solution when the free transport equation is used to describe the thermal transpiration [36].
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=F_{in}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ F(x,v)&=m_{T(x)}\int_{u\cdot n(x)>0}F(x,u)u\cdot n(x)du,&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (1.5)
@@
$\mathbf{B}(|v-u|,\omega)=\mathfrak{B}|v-u|^{\gamma}\left(\frac{u-v}{|u-v|}\cdot\omega\right),\quad-3<\gamma\leq 1,$ (1.6)
@@
$\partial\Omega_{1}=\Big(\partial\Omega\cap\{z<a\}\Big)\cup\Big(\partial\Omega\cap\{z>b\}\Big).$ (1.7)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{5/2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (1.8)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=-\frac{1}{\kappa}L(f)+\frac{1}{\kappa}\Gamma(f,f),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f(x,v)&=f_{in}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f(x,v)&=m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (1.9)
@@
The operator $L$ has two components: the multiplicative operator $\nu$ and integral operator $K$, that is , $L(f)=\nu(v)f-K(f)$. Their definition can be found in Chapter 2 or [17]. In what follows, we discuss the well-posedness of the boundary value problem (1.9). The integral form of (1.9) is
$f=\frac{1}{\kappa}S^{\kappa}_{\Omega}Kf+J^{\kappa}\left(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+\left(m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\right)\textbf{1}_{\partial\Omega^{-}_{2}}\right)+\frac{1}{\kappa}S^{\kappa}_{\Omega}\Gamma(f,f).$ (1.10)
@@
$S_{\Omega}^{\kappa}f(x,v)$ $\coloneq\int_{0}^{\tau_{-}(x,v)}e^{-\frac{1}{\kappa}\nu(v)s}f(x-sv,v)ds,$ (1.11)
$J^{\kappa}f$ $\coloneq e^{-\frac{1}{\kappa}\nu(v)\tau_{-}(x,v)}f(q(x,v),v),$ (1.12)
@@
$\tau_{-}(x,v)=\sup\{t\,|\,x-tv\in\overline{\Omega}\}\quad\text{and}\quad q(x,v)=x-\tau_{-}(x,v)v.$ (1.13)
@@
$\langle f,g\rangle$ $=\int_{\Omega\times\mathbb{R}^{3}}fgdxdv,$ (1.14)
$\langle f,g\rangle_{\partial^{-}_{2}}$ $=\int_{\partial\Omega^{-}_{2}}fg|v\cdot n(x)|dxdv.$ (1.15)
@@
We now introduce the concept of an unsealed set $\Omega$ in $\mathbb{R}^{3}$. The term unseal describes the visibility of $\partial\Omega_{1}$ from points on $\partial\Omega$. If every point $x$ can directly see the $\partial\Omega_{1}$, then we call $\Omega$ a completely unsealed set. If, for every $x\in\partial\Omega$, the $\partial\Omega_{1}$ becomes visible after tracing back $m$ times for some fixed integer $m$, then we call $\Omega$ an unsealed set. The precise definition is in Chapter 4.
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Suppose that $\Omega$ is unsealed with respect to $\partial\Omega_{1}$ and satisfies the assumption 1.1. Assume that $3<\gamma\leq 1$ and that (1.6) holds. Then, for any $0<\alpha<1/4$ and $\kappa_{0}\in(0,\infty)$, there exists $\delta>0$ which depends on $\alpha$, $\Omega$, $\partial\Omega_{1}$, and $\kappa_{0}$ such that if
$\lVert m^{-1/2}\left(F_{in}-M\right)\rVert_{L^{\infty}_{a,0}(\partial\Omega^{-}_{1})}\leq\delta$ (1.16)
@@
$|T(x)-1|\leq\delta,\quad\forall x\in\Omega_{2}$ (1.17)
then the equation 1.5 has a unique mild solution for $\kappa\geq\kappa_{0}$. Furthermore,
$\lVert f\rVert_{L^{\infty}_{\alpha,0}}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}\Big(\left\lVert m^{-1/2}\left(F_{in}-M\right)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\left\lVert m^{-1/2}(m_{T(x)}-m)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\Big).$ (1.18)
@@
$u(x)\coloneq\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv,\quad U(z)\coloneq\int_{\{(x_{1},x_{2},z)\in\Omega\}}u(x)dS.$ (1.19)
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary such that $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex. Suppose for every $\kappa>\kappa_{0}$, there exists a constant $\mathcal{C}$ such that (1.8) admits a mild solution with the following estimate:
$\lVert F\rVert_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}\leq\mathcal{C},$ (1.20)
@@
$U(x_{3})\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (1.21)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (1.22)
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary such that $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex. Suppose that for every $\kappa>\kappa_{0}$, equation (1.8) admits a mild solution bounded in $L^{\infty}_{0,\beta}$ by a constant independent of $\kappa$, where $\beta>\max\{\gamma,0\}+5$. Then, the total flux $U(x_{3})$ of $F$ satisfies the following estimate:
$|U(x_{3})|=\mathcal{O}(\frac{1}{\kappa}).$ (1.23)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{i}}e^{-\frac{v^{2}}{2}},&&\text{on}\quad\partial B^{-}_{i},i=1,\cdots,n,\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\cup_{i=1}^{n}\partial B^{-}_{i}\right),\end{aligned}\right$. (1.24)
@@ 1.2. Organization of the article
We briefly outline the proof in this article. We use the traditional $L^{2}-L^{\infty}$ estimate together with the bootstrap method to construct a solution of (1.5).
In Chapter 2, we review some elementary properties of the linear operator $L$. In Chapter 3, inspired by [13], [24], we introduce a simpler and more efficient approach to obtaining the $L^{2}$ estimate. This approach is based on the operator $S_{\Omega}^{\kappa}K$, which is compact in $L^{2}_{0,\gamma/2}$.
In Chapter 4, we use the unsealed condition to obtain the $L^{\infty}$ estimate. This concept provides several advantages, such as reducing the weight requirement in $L^{\infty}$. Moreover, this is a key tool in proving Theorem 1.2 and Theorem 1.3.
In Chapter 5, we use original Boltzmann equation to address the problem by modeling the distribution as in (1.8), with incoming data at $\Big(\partial\Omega\cap\{z<a\}\Big)$ and $\Big(\partial\Omega\cap\{z>b\}\Big)$. We cannot apply the powerful technique developed in [6]. However, we observe that the crucial mechanism for the appearance of flux lies in the different between the incoming data and the solution of the free transport model, which satisfies (1.6). By combining this observation with the concept of the unsealed set, we complete the proof of the main theorems in this article.
Chapter 2. Preliminaries
In this chapter, we review some properties of the linear operator $L$ and its components: the multiplication operator $\nu$ and the integral operator,$K$. We also present several propositions about the operator $S^{\kappa}_{\Omega}$ and $J^{\kappa}$, which are useful in Chapter 3 and Chapter 4. In addition, we recall some basic properties of $Q_{+}$ and $Q_{-}$.
We emphasize one important point: Kawagoe Daisuke [24] and others normalize the Maxwellian as $m=e^{-|v|^{2}}$. Thus the exponents differ by a factor of 2 and the constants are slightly different.

### v2

@@ 1.1. Introduction
$\frac{P_{1}}{\sqrt{T_{1}}}=\frac{P_{2}}{\sqrt{T_{2}}}$ (3)
in equilibrium. This phenomenon is called the thermomolecular pressure difference. This relation is discussed in [25], [26]. In an experiment, S. Chu Liang provided a more accurate expression of (3) in [27], with further applications given in [2], [40].
@@
$v\cdot\nabla_{x}F(x,v)=\frac{1}{\kappa}Q(F,F)(x,v),\quad\text{in}\quad\Omega\times\mathbb{R}^{3}.$ (4)
@@
The collision operator $Q$ that appears in (4) is defined as follows.
@@
$v^{\prime}=v+((u-v)\cdot\omega)\omega\quad\text{and}\quad u^{\prime}=u-((u-v)\cdot\omega)\omega.$ (5)
@@
$v^{\prime}=\frac{v+u}{2}+\frac{|v-u|}{2}\sigma\quad\text{and}\quad u^{\prime}=\frac{v+u}{2}-\frac{|v-u|}{2}\sigma,$ (6)
@@
Consequently, stationary Boltzmann equation has been extensively applied to describe the gas distribution in the domains involving thermal transpiration. However, in engineering applications, obtaining numerical solutions requires substantial computational efforts, either deterministically ([38], [42], [46]) or stochastically ([3], [33]). Some numerical analysis of thermal transpiration employ the lattice Boltzmann model [39], the BGK model ([20], [28], [30]) or the linearized Boltzmann Equation model [41]. In particular, for the linearized Boltzmann equation model, Sone, Y. and Aoki, K. have made many significant contributions ([1], [31]). Moreover, Sone provided an explicit form of the solution when the free transport equation is used to describe the thermal transpiration [36].
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=F_{in}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ F(x,v)&=m_{T(x)}\int_{u\cdot n(x)>0}F(x,u)u\cdot n(x)du,&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (7)
@@
$\mathbf{B}(|v-u|,\omega)=\mathfrak{B}|v-u|^{\gamma}\left(\frac{u-v}{|u-v|}\cdot\omega\right),\quad-3<\gamma\leq 1,$ (8)
@@
$\partial\Omega_{1}=\Big(\partial\Omega\cap\{z<a\}\Big)\cup\Big(\partial\Omega\cap\{z>b\}\Big).$ (9)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{5/2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{u\cdot n(x)>0}F(x,u)u\cdot n(x)du,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (10)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=-\frac{1}{\kappa}L(f)+\frac{1}{\kappa}\Gamma(f,f),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f(x,v)&=f_{in}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f(x,v)&=m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (11)
@@
The operator $L$ has two components: the multiplicative operator $\nu$ and integral operator $K$, that is , $L(f)=\nu(v)f-K(f)$. Their definition can be found in section 2 or [17]. In what follows, we discuss the well-posedness of the boundary value problem (11). The integral form of (11) is
$f=\frac{1}{\kappa}S^{\kappa}_{\Omega}Kf+J^{\kappa}\left(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+\left(m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\right)\textbf{1}_{\partial\Omega^{-}_{2}}\right)+\frac{1}{\kappa}S^{\kappa}_{\Omega}\Gamma(f,f).$ (12)
@@
$S_{\Omega}^{\kappa}f(x,v)$ $\coloneq\int_{0}^{\tau_{-}(x,v)}e^{-\frac{1}{\kappa}\nu(v)s}f(x-sv,v)ds,$ (13)
$J^{\kappa}f$ $\coloneq e^{-\frac{1}{\kappa}\nu(v)\tau_{-}(x,v)}f(q(x,v),v),$ (14)
@@
$\tau_{-}(x,v)=\sup\{t\,|\,x-tv\in\overline{\Omega}\}\quad\text{and}\quad q(x,v)=x-\tau_{-}(x,v)v.$ (15)
@@
$\langle f,g\rangle$ $=\int_{\Omega\times\mathbb{R}^{3}}fgdxdv,\quad\langle f,g\rangle_{v}=\int_{\mathbb{R}^{3}}fgdv,$ (16)
$\langle f,g\rangle_{\partial^{-}_{2}}$ $=\int_{\partial\Omega^{-}_{2}}fg|v\cdot n(x)|dxdv.$ (17)
@@
We now introduce the concept of an unsealed set $\Omega$ in $\mathbb{R}^{3}$. The term unseal describes the visibility of $\partial\Omega_{1}$ from points on $\partial\Omega$. If every point $x$ can directly see the $\partial\Omega_{1}$, then we call $\Omega$ a completely unsealed set. If, for every $x\in\partial\Omega$, the $\partial\Omega_{1}$ becomes visible after tracing back $m$ times for some fixed integer $m$, then we call $\Omega$ an unsealed set. The precise definition is in section 4.
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Suppose that $\Omega$ is unsealed with respect to $\partial\Omega_{1}$ and satisfies the assumption 1.1. Assume that $3<\gamma\leq 1$ and that (8) holds. Then, for any $0<\alpha<1/4$ and $\kappa_{0}\in(0,\infty)$, there exists $\delta>0$ which depends on $\alpha$, $\Omega$, $\partial\Omega_{1}$, and $\kappa_{0}$ such that if
$\lVert m^{-1/2}\left(F_{in}-M\right)\rVert_{L^{\infty}_{a,0}(\partial\Omega^{-}_{1})}\leq\delta$ (18)
@@
$|T(x)-1|\leq\delta,\quad\forall x\in\Omega_{2}$ (19)
then the equation (7) has a unique mild solution for $\kappa\geq\kappa_{0}$. Furthermore,
$\lVert f\rVert_{L^{\infty}_{\alpha,0}}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (20)
$\leq 2C_{\zeta}\max\{1,\kappa^{-1}\}\Big(\left\lVert m^{-1/2}\left(F_{in}-M\right)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\left\lVert m^{-1/2}(m_{T(x)}-m)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\Big).$
@@
$u(x)\coloneq\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv,\quad U(z)\coloneq\int_{\{(x_{1},x_{2},z)\in\Omega\}}u(x)dS.$ (21)
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary such that $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex. Suppose for every $\kappa>\kappa_{0}>0$, there exists a constant $\mathcal{C}$ such that (10) admits a mild solution with the following estimate:
$\lVert F\rVert_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}\leq\mathcal{C},$ (22)
@@
$U(x_{3})\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (23)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (24)
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary such that $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex. Suppose that for every $\kappa>\kappa_{0}$, equation (10) admits a mild solution bounded in $L^{\infty}_{0,\beta}$ by a constant independent of $\kappa$, where $\beta>\max\{\gamma,0\}+5$. Then, the total flux $U(x_{3})$ of $F$ satisfies the following estimate:
$|U(x_{3})|=\mathcal{O}(\frac{1}{\kappa}).$ (25)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{i}}e^{-\frac{v^{2}}{2}},&&\text{on}\quad\partial B^{-}_{i},i=1,\cdots,n,\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\cup_{i=1}^{n}\partial B^{-}_{i}\right),\end{aligned}\right$. (26)
@@ 1.2. Organization of the article
We briefly outline the proof in this article. We use the traditional $L^{2}-L^{\infty}$ estimate together with the bootstrap method to construct a solution of (7).
In section 2, we review some elementary properties of the linear operator $L$. In section 3, inspired by [13], [14], and [24], we introduce a simpler and more efficient approach to obtaining the $L^{2}$ estimate and we introduce the damping term $\hat{\alpha}\nu f$ rather than $\hat{\alpha}f$ to prove Theorem 3.1 for both positive and negative $\gamma$. This approach is based on the operator $S_{\Omega}^{\kappa}K$, which is compact in $L^{2}_{0,\gamma/2}$.
In section 4, we use the unsealed condition to obtain the $L^{\infty}$ estimate. This concept provides several advantages, such as reducing the weight requirement in $L^{\infty}$. Moreover, this is a key tool in proving Theorem 1.2 and Theorem 1.3.
In section 5, we use original Boltzmann equation to address the problem by modeling the distribution as in (10), with incoming data at $\Big(\partial\Omega\cap\{z<a\}\Big)$ and $\Big(\partial\Omega\cap\{z>b\}\Big)$. We cannot apply the powerful technique developed in [6]. However, we observe that the crucial mechanism for the appearance of flux lies in the different between the incoming data and the solution of the free transport model, which satisfies (8). By combining this observation with the concept of the unsealed set, we complete the proof of the main theorems in this article.
Section 2. Preliminaries
In this section, we review some properties of the linear operator $L$ and its components: the multiplicative operator $\nu$ and the integral operator, $K$. We also present several propositions about the operator $S^{\kappa}_{\Omega}$ and $J^{\kappa}$, which are useful in section 3 and section 4. In addition, we recall some basic properties of $Q_{+}$ and $Q_{-}$.
We emphasize that we linearize around Maxwellian $M=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}}$. There is another popular convention to linearize around $m=\frac{1}{\pi^{3/2}}e^{-|v|^{2}}$, for example [24], [36]. Thus, the exponents differ by a factor of 2 and the constants are different.